The table below lists the number of people by labor force classification for the country of Springfield.

______________________________________

Employed 78 million

Part-time employed 22 million

Unemployed 12 million

Not in the Labor Force 30 million

Discouraged 4 million

______________________________________

Discouraged 4 million

(1) Calculate the number in the labor force, and the unemployment rate.

(2) Calculate the number in the adult population, and the labor force participation rate.
1
Expert's answer
2020-12-21T03:30:51-0500

1) Number in the labor force

= Employed + part-time employed + unemployed

= 78 million + 22 million + 12 million

= 112 million


Unemployment rate = "\dfrac {unemployed}{labour \space force} × 100\%"


Unemployment rate = "\dfrac {12 \space million} {112 \space million} × 100\%"

Unemployment rate = 10.714286%

= 10.7%



2) Number in labour force population

= Total labour force + discouraged workers

= 112 million + 4 million

= 116 million


Labor force participation rate

= "\dfrac {labor \space force} {Working-age \space population} ×100\%"


"= \dfrac {112 \space million} {116 \space million } × 100\%"


= 96.551724%


= 96.6%